# Compressing Scene Dynamics: A Generative Approach

### v1

@@ 2.1  The Dynamics-Codec Framework
2.3  Flow-driven Decoder In previous GVC schemes, generators are usually crafted following the structures of Generative Adversarial Network [21] or Variational Auto-Encoder [22]. The resulting GVC methods can only address single scenario application with one trained model, limiting their generalizability to more diverse contents. Recently, the Diffusion Model [12] is proposed to perform generation with the denoising process and exhibits the outstanding capability of learning complex data distribution. It can be integrated with versatile conditions for tailored generative vision tasks [23]. Motion trajectory is also leveraged as the control signal for interactive image-to-video generation [24, 25, 14, 26, 27]. However, they are not yet explored in GVC framework. In this paper, a pre-trained SVD [13] and motion-adaptor [14] are utilized to ensure the robustness of motion-driven generation and its generalizability across various contents.
Specifically, the reconstructed key frame is denoted as $\hat{\textbf{I}}$, and it is first fed into a image feature extractor $\textbf{e}_{i}$,
$\textbf{y}_{i}=\textbf{e}_{i}(\hat{\textbf{I}})$ (6)
where $\textbf{y}_{i}\in\mathbb{R}^{N_{v}\times H\times W}$ denotes the image feature. Here, the image feature extractor is designed as a U-Net like structure. The decoded motion token vector $\hat{\textbf{w}}$, $\hat{\textbf{b}}$ and the image feature are processed by a dense flow generator $\textbf{g}_{f}$,
$\hat{\textbf{m}}_{f}=\textbf{g}_{f}(\textbf{y}_{i}\cdot\hat{\textbf{w}}+\hat{\textbf{b}})$ (7)
where $\hat{\textbf{m}}_{f}\in\mathbb{R}^{H\times W\times 2}$ denotes the recovered dense motion flow and $\cdot$” denotes channel-wise multiplication. Here, the flow generator is designed as a U-Net like structure. Then, following the practice in [14], the recovered dense motion is used to warp the multi-scale features of reconstructed key frame, and the features are further fused with trainable SVD encoder features. Finally, they serve as the trajectory guidance for the pre-trained diffusion decoder. In this way, massive training data can be introduced by pre-trained models without any additional bit-rate or training cost, which can consequently guarantee the visual-pleasing reconstruction of the flow-driven decoder.
@@
2.4  Optimization To bridge the gap between proposed motion token representations and pre-trained flow-driven diffusion-based generator, the optimization of Dynamics-Codec is aimed to align recovered dense motions with the original dense motion inputs of the motion adaptor. To achieve that, only flow tokenizer and dense motion recovery module are optimized, and the original dense motion are provided as supervision signals. Here, the original motion predictor of the motion adaptor is denoted as $\phi$, and L1 loss is used to measure the coordinate-wise error between ground truth flows and generated flows. The optimization objective can be written as,
$\mathcal{L}=||\hat{\textbf{m}}_{f}-\phi(\textbf{I},\textbf{P})||_{1}$ (8)
@@
3 Validations
[…]
3.1  Experimental Settings    Dataset.
Figure 3: The overview of small motion scene dynamic test set.
We collect small motion dynamic videos with various scenes and resolution of 384 $\times$ 384 from the Internet for training and evaluation. For training, total 900 videos are used to extract their ground truth dense motions between the key frame and every inter frame. For evaluation, 35 videos with diverse video contents are utilized, as shown in Figure 3, and each sequence contains 50 frame with frame rate of 25 fps.
Training Settings. We train the flow tokenizer and dense flow recovery module jointly without entropy coding. The size of image/flow features are set as 384 $\times$ 384, and the number of motion tokens in each token vector is set as 20, yielding totally 40 parameters. We implement the networks with Pytorch framework, which are then optimized by Adam optimizer with $\beta_{1}=0.5$, $\beta_{1}=0.999$ and learning rate of $10^{-4}$. We also use multi-step learning rate scheduler with $\gamma$ = 0.1 and $milestone$ = [30, 45] We train the networks on NVIDIA GeForce RTX 3090 GPU for 50 epochs with the batch size of 24.
Evaluation Settings. For Dynamics-Codec key frame compression, VVC codec with VTM 22.2 platform is used with intra mode and QP of 32, 37, 42, 47. We compare the proposed Dynamics-Codec with VVC anchor performances, which are obtained with Low-Delay-Bidirectional (LDB) configuration and QP of 37, 42, 47, 52. For evaluation metrics, we choose three perceptual measurements that are commonly used for generative contents, i.e., Learned Perceptual Image Patch Similarity (LPIPS) [28], Deep Image Structure and Texture Similarity (DISTS) [29], and Frechet Video Distance (FVD) [30]. Additionally, we use Bjøntegaard-delta-rate (BD-rate) [31] and rate-distortion (RD) curve to compare the proposed method with anchor method. To display the RD curves in an increasing manner and calculate BD-rate saving, we use “1-DISTS”, “1-LPIPS” and “5000-FVD” as the y-axis of the graphs.
[…]
@@ 3 Validations
Figure 4: Rate-distortion performance comparisons with VVC [32] in terms of DISTS, LPIPS and FVD for scene dynamics test set.
@@
Table 1: RD performance comparisons on scene dynamics test set in terms of average BD-rate savings over the VVC anchor [32].
Rate-DISTS Rate-LPIPS Rate-FVD
-35.55% -23.98% -29.91%
@@
Subjective Quality. Subjective visual quality comparisons of proposed Dynamics-Codec and VVC are shown in Figure 5. Specifically, reconstructions of test sequence 031 at 10 kbps, sequence 029 at 6 kbps and sequence 034 at 5 kbps are provided. The motion dynamics of each sequence are shown in first columns in the format of y-t space-time slices by concatenating 5-pixels-wide samples obtained from each of 50 frame. The sampled positions are indicated by red-dash lines in the first frame. The sampled dynamics present the motion patterns in sampled position and it can be seen that our method can accurately reconstruct the temporal trajectory of the objects. Then, 10 full reconstructed frames from each sequence are presented for video content comparison. It can be observed that, under ultra-low bit-rate, VVC reconstructions exist annoying blocking effects, while the Dynamics-Codec generates more visual-pleasing reconstructions with decent motion accuracy. These results demonstrate the robust generation capability of proposed motion-driven decoder for both high-quality generation and generalizability across diverse scenes.
(a) Sequence 031 at 10 kbps
(b) Sequence 029 at 6 kbps
(c) Sequence 034 at 5 kbps
Figure 5: Visual quality comparisons on test sets among VVC [32] and Dynamics-Codec (Ours) at the similar bit rate. Please refer to our project page for more demos.
4 Conclusion
@@
In this paper, we propose to exploit motion-pattern-prior instead of video-content-prior for generative video coding. By doing so, we build a novel Dynamics-Codec for scene dynamics. In particular, the compact motion prior representations are created by identifying the primary motion components and condensing them to tokens, while the robust reconstructions among various video contents are guaranteed by powerful diffusion-based motion-driven decoder. The experiment results show that our methods can achieve more than 35% BD-rate saving against VVC as well as visual-pleasing reconstructions under ultra-low bit-rate on small motion dynamics sequences, highlighting better robustness and generalizability of the motion-prior-based generative video coding scheme.

### v4

@@ 3.4. Flow-driven Generator
where n denotes the Gaussian noise added to the key frame, $Warp(\cdot)$ denotes the warping operation, and c denotes the denoising condition. The subsequently inter frame $\hat{\textbf{P}}$ generation can be completed by the pre-trained SVD decoder $\textbf{D}_{svd}$ with the trajectory guidance from c,
(9) $\hat{\textbf{P}}=\textbf{D}_{svd}(\hat{\textbf{I}}+\textbf{n},\textbf{c})$
In this way, massive training data can be introduced by pre-trained models without any additional bit-rate or training cost, which can consequently guarantee the visual-pleasing reconstruction of the flow-driven decoder.
3.5. Optimization
To bridge the gap between proposed motion token representations and pre-trained flow-driven diffusion-based generator, the optimization of Dynamics-Codec is aimed to align recovered dense motions with the original dense motion inputs of the motion adaptor. To achieve that, only flow tokenizer and dense motion recovery module are optimized, and the original dense motion are provided as supervision signals. Here, the original motion predictor of the motion adaptor is denoted as $\phi$, and L1 loss is used to measure the coordinate-wise error between ground truth flows and generated flows. The optimization objective can be written as,
@@ 3.5. Optimization
4. Validations
4.1. Experimental Settings
4.1.1. Dataset
We collect small motion dynamic videos from the Internet, with various scenes and ranging from size of 720p to 1024p. We square-crop the main areas of scenes dynamics and resize them to resolution of 384 $\times$ 384 for training and evaluation. For training, total 900 videos are used to extract their ground truth dense motions between the key frame and every inter frame. For evaluation, we include two different datasets, as shown in Figure 5, to verify the effectiveness of proposed motion-pattern-based prior and generalizability of proposed method. For the small motion dynamic videos (denoted as motion dynamic test set in the subsequent context), 40 videos with diverse video contents are utilized, and each sequence contains 50 frames with frame rate of 25 fps. We also include talking-face contents for evaluation from JVET-GFV test set (gfvc-ctc). We select the first 50 frames of all 18 sequences from Class D, which have head-and-shoulder contents and resolution of 512 $\times$ 512, and resize them to 384 $\times$ 384 to evaluate the performance of proposed method on talking-face contents.
(a) Motion dynamic test set
(b) JVET-GFV test set
Figure 5. The overview of test sets.
4.1.2. Implementation Details
We train the flow tokenizer and dense flow recovery module jointly without entropy coding. The size of image/flow features are set as 384 $\times$ 384, and the number of motion tokens in each token vector is set as 20, yielding totally 40 parameters. We implement the networks with Pytorch framework, which are then optimized by Adam optimizer with $\beta_{1}=0.5$, $\beta_{1}=0.999$ and learning rate of $10^{-4}$. We also use cosine annealing learning rate scheduler with minimum learning rate of $5\cdot 10^{-7}$. We train the networks on NVIDIA GeForce RTX 3090 GPUs for 100 epochs with the batch size of 32.
4.1.3. Evaluation Settings
For evaluation metrics, we follow the common practice in generative video coding (cttr; mttf), and choose two perceptual measurements that are commonly used for generative contents, i.e., Learned Perceptual Image Patch Similarity (LPIPS) (lpips), Deep Image Structure and Texture Similarity (DISTS) (dists). These two metrics measure the mean square error and structural similarity on feature maps extracted by VGG network and show high correlation with human perception (gfvc-survey-trans). Additionally, we also choose Frechet Video Distance (FVD) (unterthiner2019fvd) to evaluate the temporal consistency by capturing the temporal dynamics and comparing the feature distribution between the original reconstructed videos. Natural Image Quality Evaluator (NIQE) (niqe), a commonly used general purpose no-reference metric is also included to evaluate the naturalness of the decoded videos.
@@ 4.1.3. Evaluation Settings
Figure 6. Rate-distortion performance comparisons with VVC (vvc), ECM (ecm), DCVC-FM (dcvc-fm) and TPSM (tpsm) in terms of DISTS, LPIPS, FVD and NIQE for motion dynamic test set.
(a) Rate-DISTS
(b) Rate-LPIPS
(c) Rate-FVD
(d) Rate-NIQE
Figure 7. Rate-distortion performance comparisons with VVC (vvc), ECM (ecm), DCVC-FM (dcvc-fm) and TPSM (tpsm) in terms of DISTS, LPIPS, FVD and NIQE for JVET-GFV (gfvc-ctc) test set.
4.2. Compared Algorithms
To verify the effectiveness of the proposed method, we select two state-of-the-art conventional video codec VVC (vvc), ECM (ecm), one deep-learning-based video codec DCVC-FM (dcvc-fm), and one latest generative video codecs TPSM (tpsm) for comparisons. In the following, we discuss the implementation details.
  • Conventional Codec. VVC (vvc) is the latest hybrid video coding standard and ECM is the test model of next-generation video standard, which significantly improves the rate distortion performance compared with their predecessors. We adopt VTM 22.2 platform and ECM 17.1 platform with the Low-Delay-Bidirectional (LDB) configuration, where the quantization parameters (QP) are set to 37, 42, 47 and 52.
  • Deep-learning-based Video Codec. DCVC-FM (dcvc-fm) is one of the latest deep-learning-based video codec that expands the quality range and stabilizes long prediction chain with feature modulation and outperforms ECM (ecm). We implement the DCVC-FM (dcvc-fm) with the official codebase and set the quantization factors to 0, 5, 10 and 15 for evaluation.
  • Generative Video Codec. TPSM (tpsm) is one of the state-of-the-art generative codec with thin-plate spline (TPS) transform for motion estimation. Specifically, it estimates every TPS transform with 5 key-point pairs on each frame and utilizes multi-resolution occlusion masks to improve the generation quality. The key frames are compressed by VVC codec (VTM 22.2) with QP of 32, 37, 42 and 47.
  • The proposed Dynamics-Codec. For key frame compression, we implement both VVC codec with VTM 22.2 platform (denoted as “w/VVC”) and ECM 17.1 platform (denoted as “w/ECM”) with intra mode and QP of 32, 37, 42, 47.
4.3. Evaluation Results
4.3.1. Rate-Distortion Performance
[…]
@@ 4.3.1. Rate-Distortion Performance
Table 1. RD performance comparisons on scene dynamics test set in terms of average BD-rate savings over the VVC anchor (vvc).
Algorithm Rate-DISTS Rate-LPIPS Rate-FVD Rate-NIQE
ECM (ecm) -1.07% -5.32% 2.36% 0.29%
DCVC-FM (dcvc-fm) -5.73% -22.31% -3.51% 9.36%
TPSM (tpsm) -27.26% -17.56% -23.31% 18.08%
Proposed (w/VVC) -38.59% -32.54% -35.13% -34.61%
Proposed (w/ECM) -38.80% -33.89% -25.57% -37.57%
Table 2. RD performance comparisons on JVET-GFV test set in terms of average BD-rate savings over the VVC anchor (vvc).
Algorithm Rate-DISTS Rate-LPIPS Rate-FVD Rate-NIQE
ECM (ecm) -7.00% -7.79% -4.13% -3.50%
DCVC-FM (dcvc-fm) -13.77% -31.03% -12.62% -35.18%
TPSM (tpsm) -24.57% -15.74% -27.215% -1.44%
Proposed (w/VVC) -44.91% -29.57% -34.44% -46.26%
Proposed (w/ECM) -48.50% -32.38% -44.02% -48.38%
(a) Sequence of motion dynamics test set at 6kbps
(b) Sequence of JVET-GFV test set at 10kbps
Figure 8. Subective quality comparisons with VVC (vvc), ECM (ecm), DCVC-FM (dcvc-fm) and TPSM (tpsm). Please refer to the project page for video demos.
4.3.2. Subjective Quality
@@
Subjective visual quality comparisons of proposed Dynamics-Codec and comparison methods are shown in Figure 8. Specifically, reconstructions of motion dynamics sequence at 6kbps and JVET-GFV (gfvc-ctc) sequence at 10kbps are provided. Two reconstructed frames from each sequence are presented for subjective comparison. It can be observed that, under ultra-low bit-rate, VVC (vvc) and ECM (ecm) reconstructions exist annoying blocking effects, DCVC-FM (dcvc-fm) reconstructions are blurry, and TPSM (tpsm) reconstructions show obvious deformation with larger objects movements, while the Dynamics-Codec generates more visual-pleasing reconstructions with decent motion accuracy. These results demonstrate the robust generation capability of proposed motion-driven decoder for both high-quality generation and generalizability across diverse scenes.
Furthermore, we conduct user study to compare our Dynamic-Codec with all other algorithms at similar coding bit-rate. Specifically, we choose 15 sequences and implement “two alternatives, force choice” (2AFC) subjective test with 10 participants. During the test, these selected sequences from our method and other compared algorithms are sequentially displayed in a pair-wise manner, and the participants are asked to choose one video from each pair with better quality. To avoid experimental. As shown in Table 3, these participants are more inclined to choose our reconstructed videos as the preferred video compared to other reconstruction results. In particular, our proposed method shows absolute advantage with more than 90% preference ratio with VVC (vvc), ECM (ecm) and TPSM (tpsm). As for the user preference between DCVC-FM (dcvc-fm) and ours, our reconstructed videos can still be voted with a higher ratio of 83.33%. In addition, we also provide average objective results of these tested sequences. At similar ultra-low bit-rate, our method can achieve advantageous objective quality compared with other methods.
Table 3. User preference in pairwise comparison in terms of similar coding bits consumption.
Comparisons kbps DISTS ($\downarrow$) LPIPS ($\downarrow$) FVD ($\downarrow$) NIQE ($\downarrow$) User Preference
VVC (vvc) / Ours 8.04 / 8.03 0.21 / 0.14 0.34 / 0.28 1497.99 / 546.90 9.67 / 8.68 8.00% / 92.00%
ECM (ecm) / Ours 8.15 / 8.03 0.21 / 0.14 0.33 / 0.28 1349.76 / 546.90 9.27 / 8.68 1.33% / 98.67%
DCVC-FM (dcvc-fm) / Ours 8.11 / 8.03 0.19 / 0.14 0.28 / 0.28 1052.31 / 546.90 10.22 / 8.68 16.67% / 83.33%
TPSM (tpsm) / Ours 8.28 / 8.03 0.16 / 0.14 0.30 / 0.28 960.72 / 546.90 11.79 / 8.68 6.00% / 94.00%
4.3.3. Pixel-level Quality
(a) Rate-PSNR of motion dynamic test set
(b) Rate-SSIM of motion dynamic test set
(c) Rate-PSNR of JVET-GFV test set
(d) Rate-SSIM of JVET-GFV test set
Figure 9. Pixel-level metrics comparisons with VVC (vvc), ECM (ecm), DCVC-FM (dcvc-fm) and TPSM (tpsm) in terms of PSNR, SSIM.
The pixel-level quality comparisons in terms of Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) (ssim) on two test sets are shown in Fig. 9. It can be seen that the proposed Dynamics-Codec performs inferior to VVC (vvc), ECM (ecm) and DCVC-FM (dcvc-fm) among all test sets, which is mainly due to the generative nature of the proposed method. On the one hand, the generative codecs usually sacrifice pixel-level fidelity for better perceptual quality, which is also verified in previous GVC methods (cttr; mttf; ifvc). On the other hand, pixel-level metrics shows lower correlation with human perception under low bit-rate (gfvc-survey-trans). Nevertheless, the proposed Dynamics-Codec can still achieve outperform TPSM (tpsm) in terms of Rate-PSNR and Rate-SSIM on all test sets, which further demonstrates the powerful generation ability of proposed diffusion-based generator.
Table 4. Complexity Comparison in terms of encoding time and decoding time (vvc).
Algorithm Encoding time (s) Decoding time (s) Total time (s)
VVC (ecm) 1648.16 0.73 1648.89
ECM (ecm) 143.73 0.19 143.91
DCVC-FM (dcvc-fm) 2.75 2.60 5.35
TPSM (tpsm) 20.69 13.81 34.57
Proposed (w/VVC) 26.73 117.22 143.95
Proposed (w/ECM) 203.41 119.93 323.34
4.4. Complexity Analysis
We measure the encoding time and decoding time of all methods for complexity comparisons. The experiments are conducted with Intel Xeon Silver 4210 CPU @ 2.20GHz and NVIDIA GeForce RTX 3090 GPU. The average time durations to encode and decode four sequences with all different qualities are reported in Table 4. It can be seen that, the proposed Dynamics-Codec shows competitive encoding time compared with other generative and conventional methods, while the decoding time is relatively high due to the iterative sampling process of diffusion model. In total, the proposed Dynamics-Codec shows largely reduced total time consumption compared to ECM (ecm) and VVC (vvc). Meanwhile, our methods shows higher encoding time when using ECM (ecm) as the key frame encoder.
5. Conclusion
@@
In this paper, we propose to exploit motion-pattern-prior instead of video-content-prior for generative coding of scene dynamics. With proposed novel Dynamics-Codec, compact motion prior representations are extracted by identifying primary motion components and condensing them into tokens. Meanwhile, robust reconstructions across diverse video contents are ensured by a powerful diffusion-based, motion-driven decoder. The experiment results show that our methods can achieve more than 48% BD-rate saving against VVC as well as visual-pleasing reconstructions under ultra-low bit-rate on diverse scenes, highlighting better robustness and generalizability of the motion-prior-based generative video coding scheme.